### v1

# Simultaneous Masking, Not Prompting Optimization: A Paradigm Shift in Fine-tuning LLMs for Simultaneous Translation

Abstract
Large language models (LLMs) have achieved state-of-the-art performance in various language processing tasks, motivating their adoption in simultaneous translation. Current fine-tuning methods to adapt LLMs for simultaneous translation focus on prompting optimization strategies using either data augmentation or prompt structure modifications. However, these methods suffer from several issues, such as an unnecessarily expanded training set, computational inefficiency from dumping the KV cache, increased prompt sizes, or restriction to a single decision policy. To eliminate these issues, we propose a new paradigm in fine-tuning LLMs for simultaneous translation, called SimulMask. It utilizes a novel attention mask technique that models simultaneous translation during fine-tuning by masking attention connections under a desired decision policy. Applying the proposed SimulMask on a Falcon LLM for the IWSLT 2017 dataset, we have observed a significant translation quality improvement compared to state-of-the-art prompting optimization strategies on three language pairs when averaged across four different latency regimes while reducing the computational cost.
Simultaneous Masking, Not Prompting Optimization:
A Paradigm Shift in Fine-tuning LLMs for Simultaneous Translation
Matthew Raffel     Victor Agostinelli     Lizhong Chen Oregon State University {raffelm, agostinv, chenliz}@oregonstate.edu
1 Introduction
Simultaneous translation refers to the process of producing a target output translation concurrently with an oncoming source input. In our increasingly interconnected world, where communication across languages in real-time is desired, simultaneous translation is becoming a requirement. Unfortunately, the task of simultaneous translation is extremely straining for human interpreters, and among the limited number of simultaneous human interpreters, their maximum translation length is around thirty minutes Moser-Mercer et al. (1998). As such, there is a need for machine learning models to alleviate the burden placed on simultaneous human interpreters.
Current literature has primarily focused on adapting end-to-end Transformer models Vaswani et al. (2017) to overcome the difficulties of simultaneous translation due to their reduced parameter counts and greater inference speedMa et al. (2020b). However, the recent successes of large language models (LLMs) Touvron et al. (2023); Jiang et al. (2023); Almazrouei et al. (2023) has prompted preliminary research applying them to simultaneous translation through fine-tuning and inference techniques Agostinelli et al. (2023); Wang et al. (2023); Koshkin et al. (2024); Wang et al. (2024); Guo et al. (2024). Unfortunately, most modern works have neglected the computational increases created by dumping the target sequence’s key and value (KV) cache Wang et al. (2024). Furthermore, there has yet to be a universal approach to fine-tuning LLMs for simultaneous translation that is not unnecessarily computationally expensive by either expanding the dataset through data augmentation, a process referred to as prefix fine-tuning Agostinelli et al. (2023); Wang et al. (2023); Koshkin et al. (2024) or increasing the prompt length through prompt restructuring Koshkin et al. (2024); Wang et al. (2024).
The lack of an efficient fine-tuning strategy of LLMs for simultaneous translation has led us to propose a new paradigm, referred to as SimulMask. SimulMask is a novel attention mask to model simultaneous translation during fine-tuning by redistributing the attention connections under a decision policy. By design, SimulMask is broadly applicable to both flexible and fixed decision policies, creating a path forward for future work to build upon it. Furthermore, if we eliminate injecting positional information into the keys and values through a modified ALiBi Press et al. (2021), SimulMask allows for key and value caching during simultaneous translation without accuracy degradation.
To validate the efficacy of SimulMask, we fine-tuned and evaluated 1.3 billion parameter Falcon models pre-trained on the RefinedWeb dataset Almazrouei et al. (2023); Penedo et al. (2023) using the IWSLT 2017 dataset Cettolo et al. (2017). All models were fine-tuned using a wait-k decision policy. From the averaged evaluation results across all wait-k values, we found an LLM fine-tuned with SimulMask and evaluated with KV caching outperforms an LLM fine-tuned with prefix fine-tuning and evaluated without KV caching by 7.34, 3.00, and 2.98 BLEU on the English-French, English-Dutch, and English-Italian language pairs, respectively. Furthermore, on average, with the wait-3 policy, recomputing the KV cache contributes toward 26.8 % of the computation time and 87.9 % of the FLOPs to generate a target translation on an A40 GPU, substantiating the need for SimulMask, which enables inference time KV caching to avoid those computations.
In summary, the main contributions of the paper include:
  1. 1.
    Providing insights on the shortcomings of current literature in adapting LLMs to simultaneous translation.
  2. 2.
    Proposing a novel attention masking approach to fine-tune LLMs for simultaneous translation that allows for efficient training and inference.
  3. 3.
    Demonstrating the efficacy of our proposed approach in terms of translation quality and computational costs by evaluating them on multiple language pairs across many wait-k values.
2 Background and Related Work
2.1 Masked Transformer Self-Attention
We briefly review self-attention functionality in transformers (Vaswani et al., 2017) focusing on masking behavior in Equation 1, which describes per-head attention calculations for some attention head $h$. For some input ${\bm{X}}\in\mathbb{R}^{N\times d_{model}}$ with $N$ being the sequence length and $d_{model}$ representing model dimensionality, the query matrix ${\bm{Q}}^{(h)}={\bm{X}}{\bm{W}}^{(h)}_{q}$, the key matrix ${\bm{K}}^{(h)}={\bm{X}}{\bm{W}}^{(h)}_{k}$, and the value matrix ${\bm{V}}^{(h)}={\bm{X}}{\bm{W}}^{(h)}_{v}$. The aforementioned weight matrices for some attention head $h$ are defined as ${\bm{W}}^{(h)}_{q}\in\mathbb{R}^{d_{model}\times d_{head}}$, ${\bm{W}}^{(h)}_{k}\in\mathbb{R}^{d_{model}\times d_{head}}$ and ${\bm{W}}^{(h)}_{v}\in\mathbb{R}^{d_{model}\times d_{head}}$, with $d_{head}$ representing per-attention head dimensionality. We define each individual query, key, and value in their respective matrices as ${\bm{q}}_{i}^{(h)}\in{\bm{Q}}^{(h)}$, ${\bm{k}}_{i}^{(h)}\in{\bm{K}}^{(h)}$, and ${\bm{v}}_{i}^{(h)}\in{\bm{V}}^{(h)}$. ${\bm{M}}$ is defined as an optional attention mask.
${\bm{A}}^{(h)}=\texttt{softmax}\left(\frac{{\bm{Q}}^{(h)}{{\bm{K}}^{(h)}}^{T}+{\bm{M}}}{\sqrt{d_{head}}}\right){\bm{V}}^{(h)}$ (1)
Critically, ${\bm{M}}$ functions by modeling context limitations or time-based dependencies that might exist during inference but do not exist during training/fine-tuning. For generative transformer blocks or architectures, ${\bm{M}}$ is defined as a causal attention mask, where each entry, ${\bm{M}}_{ij}$ is represented by Equation 2.
${\bm{M}}_{ij}=\begin{cases}0,&\text{if }j\leq i\\ -\infty,&\text{otherwise}\end{cases}$ (2)
2.2 Simultaneous Translation
Simultaneous translation is dictated by read-write decision policies, whereby a model will wait a specific amount of time before alternating between reading and writing in fixed or flexible intervals. One fixed decision policy for simultaneous translation that is broadly adopted as a common baseline to build on due to its effectiveness and simplicity is the wait-k policy Ma et al. (2019). As the name suggests, the wait-k policy will wait for k words before alternating between writing and reading a word. Although effective in simultaneous translation, alternative adaptive policies have gained traction, which base reading and writing on an auxiliary model or a predefined set of rules Cho and Esipova (2016); Gu et al. (2017); Zheng et al. (2019). While capable of impressive results, such adaptive policies often incur additional computational costs and delays.
A Transformer is trained for a simultaneous translation decision policy by masking attention scores in the encoder self-attention and the decoder cross-attention. In the case of the encoder self-attention, each source token is prevented from attending to future source tokens following a decision policy Ma et al. (2019). For example, if the source sequence length was 5 tokens and the first read step reads 2 tokens, the second read step reads 1 token, and the third read step reads 2 tokens, then the respective attention mask is provided in Figure 1.
Figure 1: An attention mask to model simultaneous translation for a transformer encoder during training.
Alternatively, in decoder cross-attention, each target token is prevented from attending to future source hidden states following the decision policy Papi et al. (2022a). Equation 3 expresses each entry of the decoder cross-attention mask, ${\bm{M}}_{tj}$.
${\bm{M}}_{tj}=\begin{cases}0,&\text{if }j\leq f(t)\\ -\infty,&\text{otherwise}\end{cases}$ (3)
In Equation 3, $f(t)$, is a decision policy function that denotes the cumulative number of source hidden states to read when predicting target token $t$. Each source hidden state aligns with one source input token.
2.3 Applying Large Language Models for Simultaneous Translation
LLMs have demonstrated remarkable performance on the task of neural machine translation Moslem et al. (2023); Vilar et al. (2023); Xu et al. (2023); Zhang et al. (2023); Iyer et al. (2023). Such successes have prompted recent works to extend the reach of LLMs into the realm of simultaneous translation Agostinelli et al. (2023); Wang et al. (2023); Koshkin et al. (2024); Wang et al. (2024); Guo et al. (2024). LLMs are especially promising for the field of simultaneous translation due to their strong understanding of language semantics and meaning. Intuitively, SimulMT LLMs inject holistic linguistic knowledge that could allow for correct translation decisions when facing difficult contextual obstacles (e.g., translating a verb in a target language without access to that verb in the source language).
Unfortunately, Equation 3 is no longer effective in modeling simultaneous translation for decoder-only LLMs as with the decoder of a Transformer. The reason is Equation 3 is constructed specifically for the cross-attention calculation between keys exclusively from the source and queries exclusively from the target, as in Transformers. Since LLMs perform a self-attention, including the prompt, the source, and the target, Equation 3 cannot properly mask the source from the target with the additional prompt and target sequences included in the keys, and the additional prompt and source sequences included in the queries. Furthermore, Equation 3 does not enforce the autoregressive language modeling behavior of LLMs. As such, alternative means to model simultaneous translation have been proposed, leveraging prompting optimization.
3 Prompting Optimization Methods
Current methods of fine-tuning LLMs for simultaneous translation fall into the general category of prompting optimization. We define prompting optimization as either employing data augmentation to help with prompting Koshkin et al. (2024); Wang et al. (2023); Agostinelli et al. (2023) or redefining the prompt structure Wang et al. (2024); Koshkin et al. (2024) to somewhat simulate simultaneous translation. In this section, we will cover these prompting optimization strategies in depth.
3.1 Data Augmentation
Prompting optimization focusing on data augmentation resorts to subdividing each sentence in a dataset into multiple partial sentence pairs. These partial sentence pairs mimic simultaneous translation, as simultaneous translation produces outputs with a partial input. We label such a method as prefix fine-tuning, and although the high-level procedure is identical amongst current works, the algorithms employed to obtain these partial sentence pairs are unique. In the case of Agostinelli et al. (2023), each source-target sentence pair is subdivided according to the wait-k policy such that if we order the new samples from smallest to largest, each subsequent sentence pair will have one additional target word and source word so long as the end of the target or source is not reached. Upon completion there will be $\texttt{max}(|S|-(k-1),|T|)$ sentence pairs, where $|S|$ and $|T|$ are the original source and target sequence lengths. The approach requires the model to predict only the final target word in the sequence during fine-tuning.
Alternatively, Wang et al. (2023) randomly sampled a subset of sentence pairs from the dataset and truncated the source sentence to be 20% to 80% of the full length according to a uniform distribution. They obtained the respective target translations by prompting ChatGPT. The new truncated source-target sentence pairs were then added to the complete dataset to expand the overall fine-tuning dataset.
3.2 Prompt Restructuring
[…]
Similarly, Koshkin et al. (2024) leverages prompt restructuring; however, it also employs prefix finetuning. Like the conversational prompting structure, it constructs a fine-tuning prompt by aligning words between the source and target sequence to mimic an oracle decision policy. However, it deviates from conversational prompting by ensuring the alignment using padding tokens in the target sequence. Then, the causally aligned sentence prompt is subdivided using a prefix fine-tuning strategy to expand the dataset with partially filled source-target sentence pairs. At inference, the LLM contains the decision policy outputting padding tokens whenever it requires more source context tokens.
4 Analysis and Shortcomings of Prompting Optimization Methods
Prompting optimization, while works to a certain degree, is inherently deficient, possessing a host of fine-tuning and inference issues. These issues include a persistent fine-tuning-inference mismatch, consistent positional confusion in the target sequence, and high computational costs. We will first present the fine-tuning-inference mismatch, followed by the positional confusion problem, and finish with the computational burdens facing the current paradigm.
4.1 Fine-tuning/Inference Mismatch
A fine-tuning-inference mismatch is a mismatch between an LLM’s fine-tuning and inference environments. For instance, fine-tuning an LLM for neural machine translation where the entire sentence is available and deploying it for simultaneous translation where little of the sentence is available when beginning generation will create a massive inference time fine-tuning-inference mismatch. Furthermore, the LLM must be fine-tuned to accommodate KV caching, the process of caching the keys and values at inference to prevent recomputation. Overall, fine-tuning for simultaneous translation aims to minimize the fine-tuning-inference mismatch.
Unfortunately, prefix fine-tuning precludes high-quality simultaneous translation with KV caching Agostinelli et al. (2023); Koshkin et al. (2024); Wang et al. (2023) as with the continuously increasing prompt size, each subsequent entry in the KV cache continuously becomes outdated, a scenario the LLM was not fine-tuned to handle. For example, suppose we have an LLM in the middle of simultaneous translation using KV caching adhering to a wait-1 policy with the following prompting structure: “Translate the following sentence from English to German: $s_{1},s_{2},...,s_{i}$ [a]: $t_{1},t_{2},...,t_{i}$”. Now suppose we read $s_{i+1}$. Then, in the subsequent write decision, the LLM must produce $t_{i+1}$. At the given write step, the KV cache entries for $t_{1},t_{2},...,t_{i}$ are each generated with a different subset of the source sequence $s_{1},s_{2},...,s_{i}$. For instance, the keys and values that generated $t_{1}$ could only attend to $s_{1}$, whereas the keys and values of $t_{i}$ could attend to $s_{1},s_{2},...,s_{i}$. However, during prefix fine-tuning, the LLM is fine-tuned to predict $t_{i+1}$ as if the KV cache for $t_{1},t_{2},...,t_{i}$ are each generated with the same subset of the source sequence $s_{1},s_{2},...,s_{i}$. As such, the KV cache for $t_{1}$ and $t_{i}$ would both be generated by attending to the source sequence $s_{1},s_{2},...,s_{i}$. Such fine-tuning-inference mismatch is unsolved through conventional prompting structures and is necessitated to increase translation quality.
Prompting restructuring also creates additional fine-tuning-inference mismatches. In Koshkin et al. (2024); Wang et al. (2024), they all fine-tune for an oracle decision policy. However, at inference, such an oracle decision policy is not truly achievable, creating a mismatch. Furthermore, since the LLMs that leverage prompt restructuring encapsulate a specific oracle decision policy into their fine-tuning curriculum, extending them to alternative decision policies at inference is infeasible without incurring a mismatch. As such, there is a need for a flexible method adaptable to a range of decision policies that also eliminates the fine-tuning-inference mismatch.
4.2 Positional Confusion
Positional confusion describes the process whereby the relative and/or global positional information during simultaneous translation progressively becomes incorrect. Unfortunately, most simultaneous translation LLMs using KV caching suffer from this positional confusion Agostinelli et al. (2023); Koshkin et al. (2024); Wang et al. (2023). The reason is that as the source sequence grows with simultaneous translation, the target sequence shifts, necessitating the target sequence’s positional information to follow suit. However, since KV caching is employed, updating the keys and values is infeasible, causing them to hold incorrect positional information.
Aligning with our previous example, the sequence portion that would experience this positional confusion would be “[a]: $t_{1},t_{2},...,t_{i}$”. For example, after the first prediction step, the positional distance between $s_{1}$ and [a]: and $t_{1}$ would be 1 and 2, respectively. Then, after the second read, where the source sequence is $s_{1},s_{2}$ the positional distance between $s_{1}$ and [a]: and $t_{1}$ would change to 2 and 3, respectively. However, while using KV caching, this positional distance would remain 1 and 2 in the keys and/or values for subsequent predictions, causing positional confusion. Continuing translation would see the gap between the true positional distance and the positional distance in the KV cache grow. Identifying an effective method to deal with positional confusion is a requirement to prevent LLM hallucinations.
4.3 Computational Inefficiency
Avoiding KV caching and recomputing all the keys and values at each prediction step is the default solution for resolving the aforementioned fine-tuning-inference mismatch and positional confusion problems while employing prefix fine-tuning. Although effective from a translation quality standpoint, doing so incurs a large computational cost, an undesirable result for streaming tasks like simultaneous translation where latency is equally important.
Outside of KV caching, the computational costs necessary for prefix fine-tuning methods are excessive. By subdividing each sample into multiple, the dataset drastically expands, contributing toward an increased cost to complete each epoch Agostinelli et al. (2023); Koshkin et al. (2024); Wang et al. (2023). Such an increase causes the duration of each epoch to rise by upwards of a factor of 5. However, unlike normal methods of expanding a dataset through data augmentation, prefix fine-tuning does not add additional information. It is from this added computational burden that Agostinelli et al. (2023); Wang et al. (2023) are forced to fine-finetune with a subset of their entire prefix datasets.
Alternatively, methods of restructuring the prompt as in Koshkin et al. (2024); Wang et al. (2024) have computational burdens of their own. For instance, Wang et al. (2024) requires adding delimiting tokens in the prompt sequence, expanding the sequence length. Similarly, the requirement of padding tokens to induce a causal alignment between the source and target sequences, as in Koshkin et al. (2024), also expands the sequence length. Since the computational cost of the self-attention cost in the LLM scales quadratically with the sequence length, such a method is undesirable for both inference and fine-tuning.
Currently, no computationally efficient fine-tuning approach exists that enables computationally efficient inference. Identifying such a method is necessitated by the desire for low latency and high-quality simultaneous translations and reducing the already high computational costs of fine-tuning LLMs.
(a) The attention for the first prediction step.
(b) The inference mirrored attention for the first prediction step.
Figure 2: The matching attention connections during inference and finetuning for simultaneous translation.
5 SimulMask: A Paradigm Shift
In this work, we propose SimulMask, a paradigm shift in fine-tuning LLMs for simultaneous translation that eschews current methods of prompting optimization. It is through SimulMask, a method for restricting attention connections during fine-tuning, that we efficiently solve the fine-tuning-inference mismatch and positional confusion problem. Although we specifically apply SimulMask for the wait-k decision policy, it is broadly applicable to a range of decision policies. In this section, we will (1) explain the process of restricting attention connections to model simultaneous translation during fine-tuning, (2) describe SimulMask, a masking method for inducing these restricted attention connections, and (3) present a solution for positional confusion that enabled by SimulMask.
5.1 Inference Mirrored Attention
Under simultaneous translation, the latest translation token at each prediction step is conditioned only on the running source tokens. Specialized attention masks on the Transformer could achieve such conditioning; however, directly mapping these to LLMs is impossible since they fail to enforce autoregressive language modeling and cannot mask properly when the prompt, source, and target sequences are collectively included in the queries and keys. As such, prior works attempted to achieve such conditioning during fine-tuning using prompting optimization strategies littered with shortcomings. We aim to return to a modeling simultaneous translation with attention masks by creating inference mirrored attention. Through inference mirrored attention, we mirror the attention connections during inference at fine-tuning according to the chosen decision policy.
As an example, suppose we model the attention connections for a wait-1 decision policy where the complete oracle input sequence is “$p_{1},s_{1},s_{2},s_{3},s_{4},p_{2},t_{1},t_{2},t_{3},t_{4}$”. In the sequence, $s_{1},s_{2},s_{3},s_{4}$ and $t_{1},t_{2},t_{3},t_{4}$ are the 4-word source and target sequences and $p_{1}$ and $p_{2}$ are prompting regions. Then, at inference, by definition of the wait-1 policy, $p_{2}$ predicts $\hat{t}_{1}$ while conditioned on the partial sequence $p_{1},s_{1},p_{2}$. As such, as shown in Figure 2(a) the query of $p_{2}$ attends to the keys of $p_{1},s_{1},p_{2}$. Thus, during fine-tuning, to eliminate the fine-tuning-inference mismatch, the query of $p_{2}$ is limited to similarly attend to the keys of $p_{1},s_{1},p_{2}$ as shown in Figure 2(b) rather than the entire source sequence. For each successive prediction step, the previously predicted target word, $t_{i}$, predicts the next target word, $t_{i+1}$ by conditioning on an extra source word, $s_{i}$, acquired from the previous read step. To mimic such behavior at fine-tuning, the query for $t_{i}$ attends to identical keys as its inference step. We provide the attention connections for the remainder of the queries in our example in Appendix A.
5.2 SimulMask
To achieve the above inference mirrored attention, we opt for an attention mask to restrict attention during fine-tuning to mimic an arbitrary decision policy during simultaneous translation. An attention mask is preferable to prompting optimization as it is flexible and directly extends the LLM causal attention mask. We call such an attention mask SimulMask.
As a demonstration, let us create a SimulMask for the wait-1 policy that extends our example from Section 5.1. Since the LLM is autoregressive, SimulMask begins with a causal attention mask from which attention connections are removed to be identical to attention connections during simultaneous translation. If we recall from our example, the prompt, $p_{2}$, generates the first target token, $t_{1}$ conditioned on $p_{1},s_{1},p_{2}$. As such, SimulMask eliminates the attention between $p_{2}$ and $s_{2},s_{3},s_{4}$. Similarly, $t_{1}$ and $t_{2}$ are conditioned on $p_{1},s_{1},s_{2},p_{2},t_{1}$ and $p_{1},s_{1},s_{2},s_{3},p_{2},t_{1},t_{2}$, respectively. Thus, attention connections are eliminated between $t_{1}$ and $s_{3},s_{4}$ and $t_{2}$ and $s_{4}$. The SimulMask for the aforementioned wait-1 policy is provided in Figure 3.
Figure 3: The SimulMask for modeling simultaneous translation according to a wait-1 decision policy during fine-tuning.
Since each decision policy performs read/write decisions differently and each limits attention differently, this requires a unique attention mask for every sentence. However, this can be done straightforwardly. The general procedure to construct a SimulMask for a given policy and sentence consists of the following steps:
  1. 1.
    Construct a causal attention mask using Equation 2 as a starting point for SimulMask.
  2. 2.
    Starting from the intersection between the query that predicts the first target token and the first source key, apply the sub-attention mask expressed in Equation 3. The sub-attention mask prevents the target queries from attending to source keys following the arbitrary decision policy.
  3. 3.
    Mask any non-source queries before the query predicting the first target token from attending to the source keys not included in the first read decision. Such a step is necessary to prevent the hidden states associated with these queries from holding information of the entire source sequence at later layers in the LLM.
The computation for constructing an arbitrary SimulMask is negligible compared with the computation of the LLM’s forward and backward passes during fine-tuning. Since SimulMask is not applied during inference, it does not impact computational cost at deployment. Therefore, SimulMask is a necessary option for mimicking simultaneous translation during fine-tuning and providing low-latency translations at inference.
5.3 Positional Reordering
Since positional confusion during inference is a byproduct of retaining outdated positional information in either the keys or values, bypassing it requires providing a form of positional information without injecting it directly into the sequence or KV cache. One positioning method that satisfies such a constraint is the popular ALiBi, which supplies positional information through biases in attention Press et al. (2021). The bias is applied to each query-key dot product row in the attention calculation as shown in Equation 4, where $m$ is a head-specific constant.
${\bm{q}}^{(h)}_{i}{{\bm{K}}^{(h)}}^{T}+{\bm{M}}_{i}+m\cdot[-(i-1),...,-1,0]$ (4)
Though simple, ALiBi has demonstrated an ability to extrapolate to much larger sequence lengths than other state-of-the-art positional encodings, making it desirable for LLMs.
Unfortunately, ALiBi, by default, does not mesh with SimulMask due to SimulMask removing attention connections between the target queries and source keys. These removed attention connections create a gap in ALiBi biases during fine-tuning that are not present at inference. An example of such a gap is provided in Figure 4(a). In Figure 4(a), the gap is present for $q_{4}$ and $q_{5}$.
(a) Original ALiBi.
(b) Modified ALiBi.
Figure 4: The ALiBi biases for a 5x5 attention matrix with SimulMask.
To eliminate the bias gap, we modify ALiBi by reducing the bias values of all query rows influenced by SimulMask. For each query row, the reduction in bias values is equivalent to the number of attention connections removed along the row using SimulMask. Figure 4(b) provides an example of such a modification. In the case of $q_{4}$, it is no longer able to attend to $k_{2}$ and $k_{3}$; therefore, the bias on the right of the gap is reduced by 2. In modifying SimulMask, we eliminate positional confusion from the LLM during simultaneous translation.
6 Experimental Setup
6.1 Fine-tuning
Our fine-tuning experiments used the Simul-LLM framework Agostinelli et al. (2023). Each experiment used a 1.3 billion parameter Falcon model pre-trained on the RefinedWeb dataset Penedo et al. (2023). The model architecture consisted of 24 layers. Each self-attention calculation had 32 attention heads, and the hidden size was 2048. We refer to models fine-tuned with prefix fine-tuning that recompute and do not recompute the KV cache as falcon-prefix-rec and falcon-prefix-norec, respectively. The models fine-tuned with SimulMask with and without modifying ALiBi are named as falcon-simul-norec-mod falcon-simul-norec, respectively. Finally, the model fine-tuned with a causal attention mask and evaluated recomputing the KV cache is named as falcon-causal-rec. For comparison purposes, we fine-tuned a Falcon model with a causal attention mask and evaluated it for neural machine translation. We refer to it as falcon-offline.
All fine-tuning was conducted on a single H100 GPU with bfloat16 precision. The non-prefix fine-tuning duration consisted of 2 epochs using a batch size of 64, whereas the prefix fine-tuning duration consisted of 1 epoch using a batch size of 1024. We applied a learning rate of $2e^{-4}$, a weight decay of 0.1, and an inverse square root scheduler. Our optimizer was AdamW Loshchilov and Hutter (2017). The fine-tuned falcon models all used ALiBi. All fine-tuning used the English-French (en-fr), English-Italian (en-it), and English-Dutch (en-nl) language pairs of the IWSLT 2017 training set Cettolo et al. (2017).
6.2 Evaluation
We evaluated translation quality and latency for simultaneous translation using Simul-LLM inference agents Agostinelli et al. (2023) interfacing with the SimulEval toolkit Ma et al. (2020a). The translation quality was determined using detokenized BLEU with SacreBLEU Post (2018). Alternatively, the latency was determined using Length-Adaptive Average Lagging (LAAL) Papi et al. (2022b). The computational cost of simultaneous translation was recorded with FLOPs, and time in seconds to complete translation. All metrics were obtained on a single A40 GPU with bfloat16 precision. Each model was evaluated at a wait-k four lower, for which it was fine-tuned. For instance, a model fine-tuned for wait-5 was evaluated at wait-1.
7 Results
Target Length Frequency Pred Time (s) Rec Time (s) Pred GFLOPs Rec GFLOPs
0-10 5.3 % 0.40 0.07 27.60 19.87
10-20 24.5 % 0.80 0.23 46.91 95.08
20-30 24.1 % 1.31 0.44 71.79 252.77
30-40 18.4 % 1.81 0.70 96.54 524.82
40-50 8.6 % 2.39 0.93 124.62 866.00
50-60 7.5 % 2.95 1.21 151.98 1352.39
60-70 3.1 % 3.43 1.41 175.44 1839.24
70-80 2.4 % 3.94 1.42 200.67 1897.82
80-90 2.1 % 4.53 1.67 229.13 2467.18
90-100 1.5 % 5.07 1.84 255.48 2870.21
Table 1: The computation measurements using the wait-3 policy for falcon-causal-rec.
7.1 Translation Quality and Latency Results
In this section, we demonstrate the efficacy of fine-tuning with SimulMask compared to fine-tuning with a causal attention mask or with prefix fine-tuning using BLEU scores and LAAL. All evaluations are performed across wait-1, wait-3, wait-5, and wait-7 policies. The translation quality and latency results on the English-French, English-Dutch, and English-French language pairs are provided in Figures 5, 6, and 7.
Figure 5: The translation quality plotted against latency for LLMs on the English-French language pair.
Figure 6: The translation quality plotted against latency for LLMs on the English-Dutch language pair.
Figure 7: The translation quality plotted against latency for LLMs on the English-Italian language pair.
Notably, in Figures 5, 6, and 7, falcon-simul-norec-mod outperforms falcon-causal-rec across all wait-k values. On average, the increase across these wait-k values in terms of translation quality is 5.44, 8.38, and 4.78 BLEU on the English-French, English-Dutch, and English-Italian language pairs. Even more impressive, falcon-simul-norec-mod outperforms falcon-prefix-rec at low wait-k values and matches it for higher wait-k values where, on average, across all wait-k values, it increases BLEU score by 7.34, 3.00, and 2.98 BLEU on the English-French, English-Dutch, and English-Italian language pairs. Such increases in translation quality reflect the efficacy of SimulMask in adapting an LLM for simultaneous translation during fine-tuning.
Figures 5, 6, and 7 also provide an ablation demonstrating the importance of modifying ALiBi with SimulMask for high-quality translations by comparing falcon-simul-norec-mod with falcon-simul-norec. For each wait-k value and language pair, falcon-simul-norec-mod outperforms falcon-simul-norec. Quantified when averaged over all wait-k values, falcon-simul-norec-mod provides an increase of 1.20, 4.56, and 3.03 BLEU on the English-French, English-Dutch, and English-Italian language pairs over falcon-simul-norec. Unsurprisingly, at higher wait-k values where the setting approaches neural machine translation, the difference in BLEU scores becomes less pronounced between the models.
A secondary ablation is provided in Figures 5, 6, and 7 by comparing falcon-prefix-rec and falcon-prefix-norec. Doing so demonstrates translation quality increasing by recomputing the KV cache across all wait-k values. When averaged across these wait-k values, the increase is 3.58, 7.48, and 1.72 BLEU on the English-French, English-Dutch, and English-Italian language pairs. Furthermore, as with the previous ablation, the difference in the BLEU score becomes less pronounced for the higher wait-k values.
One observation that initially may seem strange is that models evaluated at lower wait-k values have their LAAL deviate from their respective k to a greater degree than those evaluated at higher wait-k. Such an increase is a byproduct of the lower wait-k models generating longer predictions than their corresponding references. The increased generation length is a byproduct of the model hallucinating on sequences provided insufficient context, causing them to repeat a single token multiple times. The most pronounced occurrence of this hallucination is provided by falcon-prefix-norec and falcon-prefix-rec on the English-French language pair. The output predictions from both these models showcase repeating single tokens or phrases causing not only increased LAAL but decreased BLEU score.
7.2 Compuational Saving Results
While not using KV caching, the computation for each write step of simultaneous translation comprises of (1) recomputing the keys and values to reflect the source sequence update and (2) generating the new target tokens. In this section, we report the average time and GFLOPs required to recompute keys and values of the predicted target sequence (Rec Time and Rec GFLOPs) and generate the new target tokens (Pred Time and Pred GFLOPs) for each translation. In doing so, we demonstrate the necessity of enabling KV caching for simultaneous translation.
All computational evaluations are performed with falcon-causal-rec using the wait-3 policy on a random sampling of 1000 samples from the English-French language pair of the IWSLT 2017 test set and are reported in Table 1. The computation measurements in Table 1 are an average of samples grouped by the number of tokens in their output translations (Target Length). The frequency that a predicted translation from falcon-causal-rec is of a specified length range is also reported (Frequency). The results for wait-1, wait-5, and wait-7 policies are provided in Appendix B.
From Table 1, in terms of GFLOPs, the recomputation of the keys and values contributes toward a greater portion of the total computation than the new target token generation. However, inversely, more time is spent generating new target tokens than recomputing the keys and values. The reason for the discrepancy is the recomputation of the keys and values is performed in parallel on the A40 GPU whereas the generation of the new target tokens is a sequential operation. On average, across all target lengths, taking into account their frequencies, the recomputation of the KV cache contributed to 26.8% of the computation time and 87.9% of the FLOPs required to generate a target translation at wait-3.
8 Conclusion
In this work, we examine current LLM fine-tuning approaches for simultaneous translation and identify their shortcomings. We then propose a new paradigm for fine-tuning LLM for simultaneous translation using a novel SimulMask that avoids the shortcomings of previous methods. Using SimulMask, the target sequence is prevented from attending to a portion of the source sequence according to an arbitrary decision policy modeling simultaneous translation. Through the application of SimulMask, we can efficiently fine-tune an LLM for simultaneous translation and reduce the computational costs of inference by eliminating the recomputation of the KV cache for the target sequence, unlike prior works. Furthermore, we can exceed or match the translation quality of prior works at all wait-k values across multiple language pairs.
Appendix A Inference Mirrored Attention Example
(a) The attention for the first prediction step.
(b) The inference mirrored attention for the first prediction step.
(c) The attention for the second prediction step.
(d) The inference mirrored attention for the second prediction step.
(e) The attention for the third prediction step.
(f) The inference mirrored attention for the third prediction step.
Figure 8: The attention during inference and finetuning for simultaneous translation.
Figure 9: A color-coded SimulMask for the wait-1 policy.
Figure 10: Color-coded inference mirrored attention produced by SimulMask for the wait-1 policy.
Appendix B Extended Computational Costs
Wait-k Length Frequency Pred Time (s) Rec Time (s) Pred GFLOPs Rec GFLOPs
1 0-10 27.9 % 0.30 0.09 22.32 27.21
1 10-20 33.0 % 0.73 0.29 43.46 125.00
1 20-30 17.2 % 1.28 0.52 70.76 315.79
1 30-40 8.5 % 1.82 0.76 96.91 592.41
1 40-50 4.7 % 2.35 0.98 123.13 920.19
1 50-60 3.2 % 2.90 1.19 149.76 1314.32
1 60-70 1.6 % 3.46 1.43 176.84 1868.62
1 70-80 1.1 % 3.99 1.55 202.72 2197.84
1 80-90 0.9 % 4.51 1.72 229.44 2725.79
1 90-100 0.6 % 5.04 1.70 254.21 2682.38
3 0-10 5.3 % 0.40 0.07 27.60 19.87
3 10-20 24.5 % 0.80 0.23 46.91 95.08
3 20-30 24.1 % 1.31 0.44 71.79 252.77
3 30-40 18.4 % 1.81 0.70 96.54 524.82
3 40-50 8.6 % 2.39 0.93 124.62 866.00
3 50-60 7.5 % 2.95 1.21 151.98 1352.39
3 60-70 3.1 % 3.43 1.41 175.44 1839.24
3 70-80 2.4 % 3.94 1.42 200.67 1897.82
3 80-90 2.1 % 4.53 1.67 229.13 2467.18
3 90-100 1.5 % 5.07 1.84 255.48 2870.21
5 0-10 4.5 % 0.35 0.01 24.98 3.21
5 10-20 20.0 % 0.80 0.11 47.13 40.15
5 20-30 24.0 % 1.31 0.35 71.70 178.58
5 30-40 17.6 % 1.82 0.58 96.96 393.19
5 40-50 11.6 % 2.36 0.84 123.17 720.18
5 50-60 6.9 % 2.92 1.09 150.63 1137.89
5 60-70 4.8 % 3.40 1.26 174.20 1527.99
5 70-80 3.3 % 4.02 1.36 204.38 1744.19
5 80-90 2.3 % 4.60 1.80 232.77 2856.33
5 90-100 1.6 % 5.07 1.88 255.55 3146.94
7 0-10 3.7 % 0.38 0 26.05 0
7 10-20 19.8 % 0.80 0.04 60.47 12.11
7 20-30 22.1 % 1.31 0.22 71.96 95.39
7 30-40 18.5 % 1.83 0.49 97.44 303.22
7 40-50 12.0 % 2.35 0.71 122.54 560.15
7 50-60 7.8 % 2.94 0.98 151.09 972.22
7 60-70 5.1 % 3.44 1.16 175.45 1311.03
7 70-80 2.9 % 4.03 1.25 204.60 1491.26
7 80-90 2.3 % 4.58 1.55 230.83 2190.85
7 90-100 1.3 % 5.07 1.80 255.25 2262.61
Table 2: The relationship between the target translation length and computational costs.